# Reliable Curation of EHR Dataset via Large Language Models under Environmental Constraints

### v1

@@ Abstract
Electronic health records (EHRs) are central to modern healthcare delivery and research; yet, many researchers lack the database expertise necessary to write complex SQL queries or generate effective visualizations, limiting efficient data use and scientific discovery. To address this barrier, we introduce CELEC, a large language model (LLM)-powered framework for automated EHR data extraction and analytics. CELEC translates natural language queries into SQL using a prompting strategy that integrates schema information, few-shot demonstrations, and chain-of-thought reasoning, which together improve accuracy and robustness. On a subset of the EHRSQL benchmark, CELEC achieves execution accuracy comparable to prior systems while maintaining low latency, cost efficiency, and strict privacy by exposing only database metadata to the LLM. CELEC also adheres to strict privacy protocols: the LLM accesses only database metadata (e.g., table and column names), while all query execution occurs securely within the institutional environment, ensuring that no patient-level data is ever transmitted to or shared with the LLM. Ablation studies confirm that each component of the SQL generation pipeline, particularly the few-shot demonstrations, plays a critical role in performance. By lowering technical barriers and enabling medical researchers to query EHR databases directly, CELEC streamlines research workflows and accelerates biomedical discovery.
@@ 1 Introduction
In practice, however, EHR systems remain challenging for researchers to use. Extracting meaningful data often requires navigating fragmented interfaces and applying technical skills such as writing structured queries or custom analytics, which most researchers lack formal training in [30, 15]. These barriers slow down scientific progress, limit exploratory analysis, and force dependence on database specialists, leaving valuable research opportunities underutilized. Tools that allow researchers to access EHR data and conduct basic analytics without programming expertise could therefore lower entry barriers and accelerate biomedical discovery [8].
Prior work has attempted to bridge this gap using text-to-SQL methods. While these approaches have shown strong performance on EHR question-answering benchmarks [10, 7, 14, 2], it is often unclear whether the underlying database systems are exposed to cloud-based LLMs—a significant risk for sensitive medical datasets, where data leakage might compromise patient privacy.
To address the security concern, we propose CELEC (Curation of EHR via LLMs under Environmental Constraints), a system that restricts LLM access to schema-level metadata only and executes all SQL queries locally, ensuring that no patient-level data are transmitted or shared inadvertently with external services. As a framework for automated EHR data extraction and analytics, CELEC translates natural language questions into SQL using a prompting strategy that combines schema information, few-shot demonstrations, and chain-of-thought reasoning to generate accurate and robust SQL outputs.
Once data are retrieved, CELEC can also generate simple visualizations directly from the extracted dataframes, providing immediate exploratory insights. Designed with real-world constraints in mind, CELEC maintains low latency and cost efficiency. CELEC also adheres to strict privacy protocols: the LLM interacts only with database metadata (e.g., table and column names), while all query execution occurs securely within the institutional environment. At no point is patient-level data transmitted to or shared with the LLM. By lowering technical barriers, CELEC empowers researchers to query EHR databases and perform preliminary analytics independently, streamlining workflows and accelerating scientific discovery.
@@ 2.1 EHR system usability & data accessibility
Despite near-universal adoption of EHRs [3, 26], effective use remains limited by fragmented information, disrupted workflows, and high cognitive load—factors linked to errors, inefficiency, and burnout among both clinicians and researchers [1, 20]. Prior research has shown that extensions such as custom dashboards and enhanced visualization of clinical trends can improve usability [19, 23]. More recently, automated visualization approaches have been explored in the medical domain [33]; however, these efforts often focused on small-scale applications and electronic medical records rather than large-scale, multi-table EHR systems. CELEC advances this line of work by integrating natural language querying with flexible analytics and visualization, providing a deployment-ready framework that directly supports research-driven use cases.
Beyond usability in care delivery, secondary use of EHR data is essential for research, quality improvement, and population health management. Yet accessibility remains a barrier: healthcare professionals often lack the training required to manipulate complex database systems [15, 30]. Some prior efforts employed visual query builders grounded in fixed SQL templates [29], but these approaches sacrifice flexibility for ease of use, limiting their applicability in production environments. CELEC addresses this limitation by leveraging LLMs to generate SQL queries directly from natural language, balancing accessibility with flexibility in data extraction and analysis.
@@ 2.2 Medical text-to-SQL
General-purpose text-to-SQL models have achieved strong results on benchmarks such as Spider [32] and WikiSQL [34] [6, 24, 18]. However, performance on these datasets does not guarantee success in clinical contexts, where data complexity and privacy requirements differ substantially from business applications. Early medical-domain systems relied on templates for text-to-SQL translation, which limited their ability to address complex real-world questions [31]. The EHRSQL shared task [16, 17] advanced the field by collecting natural language queries from hospital staff and aligning them to databases such as MIMIC-III, eICU, and MIMIC-IV demo. Leading teams on the benchmark explored methods including schema-aware models [10], ensemble prompting [7], and probability-based SQL verification [14]. CELEC employs a streamlined two-call design to strike a balance between accuracy and low latency, while extending its capabilities beyond question answering to support exploratory tasks, such as cohort selection and visualization. Quantitative comparisons are presented in Section 4.
@@ 2.3 LLMs in Healthcare
Figure 1: Overview of the CELEC system design
@@ 3 Methods
We designed the CELEC system to be an end-to-end application (see Figure 1). The system will connect to an EHR database. In our implementation, we connect to DuckDB versions of MIMIC-III [12] and MIMIC-IV [11] databases, which are large EHR databases comprising de-identified health-related data, including demographics, measurements, laboratory test results, procedures, and medications. Users can enter a natural language (NL) question as input to specify the objective of their data extraction or analysis. Then, we utilize an LLM to translate the NL question into a SQL query that correctly extracts a succinct yet necessary set of data to answer the NL question. We use prompt engineering to enhance the quality of generated SQL code (Section 3.1). Next, we connect to the EHR database and execute the generated SQL code to extract corresponding data. After the requested dataframe is retrieved, we use an LLM to create a visualization that effectively answers the NL question or enables users to better explore their data (Section 3.2).
@@ 3.1 SQL generation
Few-shot demonstrations. We include in-context few-shot demonstrations of NL questions paired with their corresponding SQL queries. Demonstrations come from two sources: (a) medical literature using MIMIC datasets, where questions were adapted from published cohort criteria and manually verified (105 demos; see Appendix C), and (b) the EHRSQL benchmark [17], from which we preprocessed the training and validation sets into 4,761 high-quality pairs (see Appendix D). These sources are combined into a single demo database; during inference, CELEC selects the top-$k$ most similar demos ($k=2$) based on the cosine similarity between the embeddings of the input question and the demo questions. Embeddings are computed using all-MiniLM-L6-v2 and indexed with ChromaDB for efficient retrieval.
@@
Error handling. After SQL generation, the query is executed locally on the target database. If execution fails (commonly due to hallucinated aliases or function names), the error message is appended to the prompt and the LLM retries generation. By default, up to two retries are allowed, which balances robustness with latency (see ablation results in Section 5). Error messages are processed only by the LLM during retries and are not exposed to users.
@@ 3.2 Visualization generation
While existing text-to-SQL systems typically return query results as raw tables, many research and clinical workflows benefit from visual summaries that facilitate pattern interpretations. To address this, CELEC includes an LLM-powered visualization module. After the SQL query is executed and a dataframe retrieved, we provide the LLM with the input NL question along with column metadata from the resulting table. The model predicts (1) an appropriate visualization type (e.g., histogram, bar chart, line plot, scatterplot) and (2) the aesthetic mappings, such as which variables should be plotted on the horizontal and vertical axes.
@@ 3.3 System integration
While Sections 3.1 and 3.2 describe SQL and visualization generation individually, their integration is key to CELEC’s usability in real-world research settings. The modules are orchestrated in a unified pipeline: SQL queries are generated and executed locally, and results can either be returned directly or forwarded to the visualization module for immediate exploratory analysis. Multi-turn refinement mechanisms further ensure that users can iterate naturally, without SQL expertise.
This integration offers system-level benefits that are not captured by either module alone. First, the streamlined design, which relies on only two LLM calls per query, reduces latency compared to ensemble or verification-heavy approaches. Second, the privacy-conscious architecture ensures that at no stage do LLMs access patient-level data; only schema metadata and aggregate specifications are processed. Finally, presenting SQL and visualization under a single framework reduces the technical barrier for researchers and clinicians alike, supporting both reproducible research and practical deployment.
@@ 4.1 Dataset adaptation
To align the benchmark with CELEC, we filtered and modified the provided train, validation, and test splits. In particular, we removed unanswerable questions where the gold SQL did not execute successfully on the MIMIC-IV demo database (see Appendix D for full details). After adaptation, we used all training and validation points as few-shot candidates for SQL prompting, 10% of the new test set (78 queries) as validation data for hyperparameter tuning, and the remaining 90% (707 queries) for evaluation. By ensuring that all test queries are executable, this adaptation yields a more reliable measure of model capability. Although not identical to the official leaderboard split, the results remain informative for comparison with published systems.
@@ 4.2 Evaluation methods
4.3 Results
Table 1 compares CELEC with leading systems from the EHRSQL-2024 leaderboard. On our adapted test set, CELEC achieves 81.05% RS(0) accuracy, abstaining in only 0.14% of cases. This result is comparable to the third-place system (ProbGate, 81.92%) and exceeds the fourth-place system (KU-DMIS, 72.07%). Unlike the leaderboard-topping LG AI Research system [10], which relied on full supervised fine-tuning, CELEC achieves this performance in a training-free setting using only few-shot prompting and two LLM calls per query.
In terms of latency, CELEC processes one NL query in an average of 6.02 to 6.11 seconds across the test set. While other systems have not disclosed inference times, CELEC’s streamlined two-call design demonstrates practical feasibility for interactive use.
@@ 4.3 Results
4.4 Discussion
These results highlight CELEC’s ability to approach state-of-the-art accuracy without training or complex pipelines, while maintaining low latency and strong privacy guarantees. Ablation studies in Section 5 further confirm that few-shot prompting and retry mechanisms are critical to this performance.
@@ 5 Ablation Study
Few-shot demonstrations. In-context demonstrations had the most significant impact of all components. Without demonstrations (zero-shot), CELEC achieved only 50.21%. Adding a single demonstration raised performance sharply to 73.97%, and using two demonstrations, which is our default, further improved it to 81.05%. This trend underscores the crucial role of in-context learning in aligning the model with the structure and semantics of medical databases, with diminishing returns evident beyond two examples.
Maximum attempts. Allowing the model to retry once after a failed SQL execution improved accuracy from 79.21% (single attempt) to 81.05% (two attempts). Error analysis indicates that many first-pass failures stemmed from superficial issues such as hallucinated aliases or function names. Providing the database error message during a second attempt often corrected these, yielding modest but consistent gains with minimal latency overhead.
@@ 6 Discussion & Conclusion
We introduced CELEC, a large language model–powered framework for automated EHR data extraction and analytics. CELEC translates natural language questions into executable SQL queries and can generate simple visualizations of query results, lowering the expertise required to access and explore EHR data. Through evaluation on the EHRSQL-2024 benchmark, CELEC achieved an RS(0) score of 81.05% on our adapted test set, comparable to strong leaderboard systems, while maintaining low latency, requiring no additional training, and adhering to a privacy-conscious design. Our ablation studies further highlighted the critical role of few-shot demonstrations, schema grounding, reasoning-optimized base models, and lightweight retry mechanisms in achieving robust SQL generation. Together, these results demonstrate that CELEC is an effective and practical approach for bridging the accessibility gap between researchers and large-scale EHR databases and has the potential to streamline workflows and accelerate discovery in healthcare research.

### v2

@@ Abstract
Electronic health records (EHRs) are central to modern healthcare delivery and research; yet, many researchers lack the database expertise necessary to write complex SQL queries or generate effective visualizations, limiting efficient data use and scientific discovery. To address this barrier, we introduce CELEC, a large language model (LLM)-powered framework for automated EHR data extraction and analytics. CELEC translates natural language queries into SQL using a prompting strategy that integrates schema information, few-shot demonstrations, and chain-of-thought reasoning, which together improve accuracy and robustness. CELEC also adheres to strict privacy protocols: the LLM accesses only database metadata (e.g., table and column names), while all query execution occurs securely within the institutional environment, ensuring that no patient-level data is ever transmitted to or shared with the LLM. On a subset of the EHRSQL benchmark, CELEC achieves execution accuracy comparable to prior systems while maintaining low latency, cost efficiency, and strict privacy by exposing only database metadata to the LLM. Ablation studies confirm that each component of the SQL generation pipeline, particularly the few-shot demonstrations, plays a critical role in performance. By lowering technical barriers and enabling medical researchers to query EHR databases directly, CELEC streamlines research workflows and accelerates biomedical discovery.
@@ 1 Introduction
In practice, however, EHR systems remain challenging for researchers to use. Extracting meaningful data often requires navigating fragmented interfaces and applying technical skills such as writing structured queries or custom analytics, which most researchers lack formal training in [33, 18]. These barriers slow scientific progress, limit exploratory analysis, and necessitate reliance on database specialists, leaving valuable research opportunities underutilized. Tools that allow researchers to access EHR data and conduct basic analytics without programming expertise could therefore lower entry barriers and accelerate biomedical discovery [10].
Prior work has attempted to bridge this gap using text-to-SQL methods. While these approaches have shown strong performance on EHR question-answering benchmarks [13, 9, 17, 2], it is often unclear whether the underlying database systems are exposed to cloud-based LLMs, which can be a significant risk for sensitive medical datasets, where data leakage might compromise patient privacy.
To address the security concern, we propose CELEC (Curation of EHR via LLMs under Environmental Constraints), a system that restricts LLM access to schema-level metadata only and executes all SQL queries locally, ensuring that no patient-level data is transmitted or shared inadvertently with external services. As a framework for automated EHR data extraction and analytics, CELEC translates natural-language (NL) questions into SQL using a prompting strategy that combines schema information, few-shot demonstrations, and chain-of-thought reasoning to generate accurate and robust SQL outputs.
Once data are retrieved, CELEC can also generate simple visualizations directly from the extracted dataframes, providing immediate exploratory insights. Designed with real-world constraints in mind, CELEC maintains low latency and cost efficiency. CELEC also adheres to strict privacy protocols: the LLM interacts only with database metadata (e.g., table and column names), while all query execution occurs securely within the institutional environment. At no point is patient-level data transmitted to or shared with the LLM. By lowering technical barriers, CELEC enables researchers to query EHR databases and perform preliminary analyses independently, thereby streamlining workflows and accelerating scientific discovery.
@@ 2.1 EHR system usability & data accessibility
Despite near-universal adoption of EHRs [3, 30], effective use remains limited by fragmented information, disrupted workflows, and high cognitive load, which are commonly linked to errors, inefficiency, and burnout among both clinicians [1, 24]. Prior research has shown that extensions such as custom dashboards and enhanced visualization of clinical trends can improve usability [22, 27]. More recently, automated visualization approaches have been explored in the medical domain [37]; however, these efforts often focused on small-scale applications and electronic medical records rather than large-scale, multi-table EHR systems.
Beyond usability in care delivery, secondary use of EHR data is essential for research, quality improvement, and population health management. Nevertheless, accessibility remains a major barrier, as healthcare professionals often lack the training required to manipulate complex database systems [18, 33]. To reduce this barrier, several interactive EHR query tools have been proposed. For example, i2b2[23] provides a drag-and-drop cohort builder in which users browse a hierarchical ontology of clinical concepts (e.g., diagnoses, laboratory tests, medications, and procedures) and place them into panels such as “Inclusion Criteria” and “Exclusion Criteria” to construct a query. Subsequent systems, such as LEAF[7] and OMOP/ATLAS[11], follow a similar paradigm, augmenting it with modern web interfaces and richer configuration options to make the experience more intuitive. While these tools eliminate the need to write SQL explicitly, they still require users to think in terms of database structure and to formulate queries using controlled vocabularies and formal logic. As a result, effective use often demands substantial familiarity with the underlying data model and query logic, which limits accessibility for users with minimal exposure to databases and programming and constrains rapid exploratory analysis.
In contrast, the goal of CELEC is to advance accessibility to the next level. CELEC removes the need for users to understand query logic, database structure, or any programming language. Instead, users simply pose a natural-language question, and the system automatically retrieves relevant data, performs the necessary analysis, and produces both tabular results and visual summaries. By unifying natural language querying with flexible analytics and visualization within a single pipeline, CELEC provides a deployment-ready framework that directly supports research-driven use cases. This question-driven interaction paradigm offers a substantially more intuitive experience than schema-aware query builders and fundamentally distinguishes CELEC from existing interactive EHR tools.
@@ 2.2 Medical text-to-SQL
General-purpose text-to-SQL models have achieved strong results on benchmarks such as Spider [36] and WikiSQL [38] [8, 28, 21]. However, performance on these datasets does not guarantee success in clinical contexts, where data complexity and privacy requirements differ substantially from business applications. Early medical-domain systems relied on templates for text-to-SQL translation, which limited their ability to address complex real-world questions [34]. The EHRSQL shared task [19, 20] advanced the field by collecting natural language queries from hospital staff and aligning them to databases such as MIMIC-III, eICU, and MIMIC-IV demo. Leading teams on the benchmark explored methods including schema-aware models [13], ensemble prompting [9], and probability-based SQL verification [17]. CELEC employs a streamlined two-call design that balances accuracy and low latency while strictly adhering to privacy protocols. Relevant quantitative evaluations are presented in Section 4.
@@ 2.3 LLMs in Healthcare
Figure 1: Overview of the CELEC system design.
@@ 3 Methods
We designed the CELEC system [^1] The source code and data will be released through GitHub. as an end-to-end application (see Figure 1). The system will connect to an EHR database. In our implementation, we connect to DuckDB versions of MIMIC-III [15] and MIMIC-IV [14] databases, which are large EHR databases comprising de-identified health-related data, including demographics, measurements, laboratory test results, procedures, and medications. Users can enter a natural-language (NL) question as input to specify the objective of their data extraction or analysis. Then, we use an LLM to translate the NL question into an SQL query that extracts a succinct yet necessary set of data to answer it. We use prompt engineering to enhance the quality of generated SQL code (Section 3.1). Next, we connect to the EHR database and execute the generated SQL to retrieve the corresponding data. After the requested dataframe is retrieved, we use an LLM to create a visualization that effectively answers the NL question or enables users to better explore their data (Section 3.2).
@@ 3.1 SQL generation
Few-shot demonstrations. We include in-context few-shot demonstrations of NL questions paired with their corresponding SQL queries. Demonstrations come from two sources: (a) medical literature using MIMIC datasets, where questions were adapted from published cohort criteria and manually verified (105 demos; see Appendix C), and (b) the EHRSQL benchmark [20], from which we preprocessed the training and validation sets into 4,761 high-quality pairs (see Appendix D). These sources are combined into a single demo database; during inference, CELEC selects the top-$k$ most similar demos ($k=2$) based on cosine similarity between the input question’s embedding and the embeddings of the demo questions. Embeddings are computed using all-MiniLM-L6-v2 and indexed with ChromaDB for efficient retrieval.
@@
Error handling. After SQL generation, the query is executed locally on the target database. If execution fails (typically due to hallucinations of aliases or function names), the error message is appended to the prompt, and the LLM retries the generation. By default, up to two retries are allowed, which balances robustness with latency (see ablation results in Section 5). Error messages are processed only by the LLM during retries and are not exposed to users.
@@ 3.2 Visualization generation
As many research and clinical workflows benefit from visual summaries that facilitate pattern interpretations, CELEC also includes an LLM-powered visualization module. After the SQL query is executed and a dataframe retrieved, we provide the LLM with the input NL question and the column metadata from the resulting table. The model predicts (1) an appropriate visualization type (e.g., histogram, bar chart, line plot, scatterplot) and (2) the aesthetic mappings, such as which variables should be plotted on the horizontal and vertical axes.
@@ 3.3 System integration
While Sections 3.1 and 3.2 describe SQL and visualization generation individually, their integration is key to CELEC’s usability in real-world research settings. The modules are orchestrated in a unified pipeline: SQL queries are generated and executed locally, and results can either be returned directly or forwarded to the visualization module for immediate exploratory analysis.
This integration offers system-level benefits that neither module alone captures. First, the streamlined design, which relies on only two LLM calls per query, reduces latency compared to ensemble or verification-heavy approaches. Second, the privacy-conscious architecture ensures that at no stage do LLMs access patient-level data; only schema metadata and aggregate specifications are processed. Finally, presenting SQL and visualization under a single framework reduces the technical barrier for researchers and clinicians alike, supporting both reproducible research and practical deployment.
@@ 4.1 Dataset adaptation
To align the benchmark with CELEC, we filtered and modified the provided train, validation, and test splits. In particular, we removed unanswerable questions where the gold SQL did not execute successfully on the MIMIC-IV demo database (see Appendix D for full details). After adaptation, we used all training and validation points as few-shot candidates for SQL prompting, 10% of the new test set (78 queries) as validation data for hyperparameter tuning, and the remaining 90% (707 queries) for evaluation. By ensuring that all test queries are executable, this adaptation yields a more reliable measure of model capability. We note that the deviation from the official leaderboard split precludes direct performance comparisons between CELEC and published systems; we present the leaderboard numbers as a contextual reference for understanding CELEC’s performance.
@@ 4.2 Evaluation methods
4.3 Results & Discussion
Table 1 compares CELEC with leading systems from the EHRSQL-2024 leaderboard. On our adapted test set, CELEC achieves 81.05% RS(0) accuracy. This result highlights CELEC’s ability to achieve accuracy comparable to existing systems while maintaining strong privacy guarantees. Ablation studies in Section 5 further confirm that few-shot prompting and retry mechanisms are critical to this performance.
In terms of latency, CELEC processes one NL query in an average of 6.02 to 6.11 seconds across the test set. As for API costs, CELEC incurs an average cost of \$0.0152 per question across the test set. While other systems have not disclosed inference times and associated costs, CELEC’s streamlined two-call design demonstrates practical feasibility for interactive use and large-scale deployment.
@@ 5 Ablation Study
Few-shot demonstrations. In-context demonstrations had the most significant impact of all components. Without demonstrations (zero-shot), CELEC achieved only 50.21%. Adding a single demonstration increased performance to 73.97%, and using two demonstrations, which is our default, further improved it to 81.05%. This trend underscores the crucial role of in-context learning in aligning the model with the structure and semantics of medical databases, with diminishing returns evident beyond two examples.
Maximum attempts. Allowing the model to retry once after a failed SQL execution improved accuracy from 79.21% (single attempt) to 81.05% (two attempts). Error analysis indicates that many first-pass failures stemmed from superficial issues, such as hallucinations of aliases or function names. Providing the database error message during a second attempt often corrected these, yielding modest but consistent gains with minimal latency overhead.
@@
6 Error Mode Analysis
To better understand the patterns underlying failed queries, we conducted an error mode analysis. We manually inspected cases where the result table produced by the predicted SQL did not match the ground-truth output and categorized the observed failures. Several recurring error modes emerged, offering insights into the model’s limitations and suggesting concrete directions for future improvement (Figure 2).
Figure 2: Statistics of failure modes under default settings.
Syntax error. Syntax errors were the most frequent error mode we observed, accounting for 35% of failures. In these cases, the predicted queries commonly included either undefined functions (e.g., “date”, which is not a DuckDB function) or invalid function argument types (e.g., for the date_diff function). A likely explanation is confusion with functions from more traditional SQL dialects, in which these queries would be valid. As these newer dialects gradually gain popularity in various applications for their expressive power and performance benefits, future work in general-domain Text-to-SQL could consider more explicit modeling of dialectal variation, potentially by developing new benchmarks.
Incorrect ranking. Incorrect ranking was the second most common error type, accounting for 27% of all failures. These errors systematically occurred for questions that requested a fixed number of ranked results, typically phrased as “what are the top K items…”. In the ground-truth SQL, such queries often rely on the DENSE_RANK function and filter results by ranks no greater than K, which allows the output to include more than K rows when ties occur at the cutoff. In contrast, the predicted SQL queries frequently use ORDER_BY in conjunction with LIMIT, which enforces an output of exactly K rows and thus fails to account for ties. This discrepancy suggests a limitation of the system in handling multiple valid interpretations of NL ranking questions. Future work could explore uncertainty-aware decoding or multi-turn interaction mechanisms that allow the system to request clarification when such ambiguities arise.
Incorrect cell values. We observed that in 10% of failures, the predicted SQL queries referenced incorrect cell values due to exact string matching, which often resulted in an empty table. Typical examples include mismatches in label specificity or casing, such as “height” versus “height (cm)” for patient height, and “M” versus “m” for male gender. As with incorrect ranking errors, these failures highlight the model’s difficulty in handling uncertainty and subtle mismatches between natural language and database-specific values. Incorporating mechanisms such as fuzzy string matching or value normalization could help improve robustness and reliability.
Incorrect dealing with multiplicity. Another observed failure mode concerns the handling of multiplicity in query answers. In some cases, the predicted SQL used a subquery that returned multiple rows in a context that expected a scalar value, which results in an invalid query. In other cases, the NL question permits multiple valid rows in the result table, but the predicted query incorrectly restricts the output to a single row.
Abstained or truncated output. The system also abstained from answering or produced truncated outputs for answerable questions, accounting for 6% of the failure cases.
Miscellaneous. Other less frequent error modes include incorrect output formats, incorrect table or column names, and arithmetic or aggregation errors.
Overall, our error analysis shows that failure cases in CELEC Text-to-SQL generation exhibit consistent and interpretable patterns. These patterns reveal limitations in the model’s ability to adapt to emergent SQL dialects and to handle ambiguity in NL queries, suggesting directions for future work.
@@ 7 Discussion & Conclusion
We introduced CELEC, a large language model–powered privacy-constrained framework for automated EHR data extraction and analytics. CELEC translates natural language questions into executable SQL queries and can generate simple visualizations of query results, lowering the expertise required to access and explore EHR data. Importantly, CELEC adheres to strict privacy protocols: the LLM accesses only database metadata and does not access patient-level data. In an evaluation on answerable MIMIC-IV cases in the EHRSQL-2024 benchmark, CELEC achieved an RS(0) score of 81.05%, comparable to existing systems, while adhering to a privacy-conscious design and maintaining low latency and low cost. Our ablation studies further highlighted the critical role of few-shot demonstrations, schema grounding, reasoning-optimized base models, and lightweight retry mechanisms in achieving robust SQL generation. Together, these results demonstrate that CELEC is an effective and practical approach in curating easily accessible EHR datasets. It has high potential to bridge the accessibility gap between researchers and large-scale EHR databases and to streamline workflows and accelerate discovery in healthcare research.
Several limitations of our work suggest directions for future research. First, our evaluation has limitations in both benchmarking and component-level assessment. To ensure executability, we filtered out unanswerable EHRSQL queries, which prevents direct comparison with systems evaluated on the original benchmark splits; additionally, we do not evaluate the visualization module and provide quantitative metrics to assess the correctness or appropriateness of the generated visualizations. Developing standardized evaluation criteria for data visualization in text-to-SQL pipelines is an important direction for future work. Second, our system is evaluated exclusively on the MIMIC database schema. While our system is not architecturally tied to MIMIC-specific table structures and terminology, further validation on additional EHR databases would help assess CELEC’s robustness and portability across different settings. Finally, we do not conduct real-world deployment evaluations. While the system is designed to lower technical barriers for clinical researchers, future user studies would help determine its practical utility, usability, and impact on research workflows.